OpenStudy (anonymous):

What is the value of f(x) = 9x when x = -2? 1 81 81 1 18 18

OpenStudy (anonymous):

1/81 81 1/18 18

OpenStudy (anonymous):

\[\huge9x // 9^x\] which one?

OpenStudy (anonymous):

first

OpenStudy (lasttccasey):

Well `f(x) = 9x` when `x=-2` so f(-2) = 9(-2) = -18 if its supposed to be `f(x) = 9^x` then f(-2) = 9^(-2) = 0.012345679

OpenStudy (anonymous):

wait no second sorry

OpenStudy (anonymous):

just plug it in. f(x) = 9x f(-2) = 9(-2) 9*-2= some number. go get a calculator

OpenStudy (anonymous):

i havent had the money to get a graphing calculator or a TI either

OpenStudy (anonymous):

start>run>calculator. google.com has one built in as does wolframalpha

OpenStudy (lasttccasey):

What he said ^